OpenStudy (anonymous):

Solve f(g(x) if:

OpenStudy (anonymous):

\[f(x)=\frac{ 1 }{ 3 }x^3-2\] g(x)=\[\sqrt[3]{3x+6}\]

OpenStudy (helder_edwin):

i think u should be able to do it on your own.

OpenStudy (helder_edwin):

\[ \large f(g(x))=f\left(\sqrt[3]{3x+6}\right)= \frac{1}{3}\left(\sqrt[3]{3x+6}\right)^3-2= \]